To test whether animal subjects consume the same amounts of sweet-tasting solutions, a researcher has subjects consume one of three sweet-tasting solutions (sucrose, saccharine, or polycose). The amount consumed (in milliliters) of each solution is given in the table. Compute eta-squared. (10 points) Type of Sweet Taste Solution 10 10 Sucrose Saccharin Polycose x2 = 2678 G 278 N 30 12 13 9 7 9 12 12 10 12 10 10 100 10 10 16 76 10 7.6 16.4 102 10 10.2 27.6
Answer #1

rectỆT 름-G2= (1002 + 762 + 1022)/10-2782/30=4186667 ss, sseffectー〉 10 30 G2 ) 30 3 10 ssrotal-Σ Σ x2-30- 2678-2782/30 STotal